At 5 pm Roger zoomed off from Asteroid. One hour later, Maryanne zoomed off from Asteroid in the opposite direction at a speed 400 kph less than that of Roger. If they were 7900 km apart at 11 p.m how fast did each travel?
Answer by stanbon(75887)

At 5 pm Roger zoomed off from Asteroid. One hour later, Maryanne zoomed off from Asteroid in the opposite direction at a speed 400 kph less than that of Roger. If they were 7900 km apart at 11 p.m how fast did each travel?
----------
Roger DATA:
Time = 6 hrs ; rate = r kph ; distance = 6r km
--------------
Maryanne DATA:
Time = 5 hrs ; rate = r-400 kph ; distance = 5(r-400) km
-------------
EQUATION:
distance + distance = 7900 km
6r + 5(r-400) = 7900
6r + 5r-2000 = 7900
11r = 9900
r = 900 kph (Roger's rate)
r-400 = 500 kph (Maryanne's rate)
===============
Cheers,
Stan H.